# Improving Out-of-Distribution Generalization of Trajectory Prediction for Autonomous Driving via Polynomial Representations

### v1

@@ I Introduction
Figure 1: Our proposed model architecture. Agent histories and road geometry are both represented via polynomials. The tracked initial states and predicted states are fused into one polynomial trajectory prediction, ensuring continuity of past observation and future prediction.
@@
For practical application, the performance of trajectory prediction should be independent of these shared biases in a single dataset and measured by Out-of-Distribution (OoD) testing, e.g., across different datasets. However, the efforts for cross-dataset evaluation are hampered by the in-homogeneous data format and task details among datasets and competitions. Notable distinctions are the different lengths of scenario, observation history, and prediction horizon in each dataset.
@@ V-D1 Augmentation Strategy
Though heterogeneous augmentation does marginally reduce FMAE’s prediction error for both ID and OoD testing, the relative and absolute increase in prediction error between ID and OoD samples is practically the same as in the non-augmented case, e.g., $+27.9$cm $(61.5\%)$ vs. $+29.5$cm $(69.7\%)$ in $\Delta minADE_{6}$.
@@
On the other hand, the robustness of QCNet significantly benefits from the homogeneous augmentation with reduced $\Delta minFDE_{1}$ from $+180.2$cm $(71.5\%)$ to $+106.7$cm $(46.4\%)$. Compared to EP-noAug, EP-F and EP-Q replicate the behavior observed from FMAE and QCNet, e.g., with only a marginal impact from heterogeneous augmentation and a significant impact from homogeneous augmentation.
@@ V-D2 Data Representation
Our EP variants exhibit improved robustness compared to benchmarks with sequence-based data independent of data augmentation strategy. For instance, EP-F outperforms FMAE by demonstrating lower prediction errors and a lower increase in prediction errors in OoD testing. Similarly, compared to QCNet, EP-Q also demonstrates improved robustness, even achieving a reduction in prediction error of $-2.2$cm $(-1.9\%)$ in $\Delta minADE_{1}$.
@@ VI Conclusion
Recent SotA trajectory prediction models have thoroughly optimized their In-Distribution (ID) performance and present outstanding results in test sets evaluated in individual prediction competition. However, better ID performance does not automatically guarantee higher Out-of-Distribution (OoD) performance. In fact, performance differences can and do reverse in OoD settings. Our proposed OoD testing protocol enables a fresh perspective for model evaluation in trajectory prediction that we hope the community will adopt. With our protocol, we demonstrate the robustness improvement from homogenous augmentation and prove the benefits of polynomial representation as employed in our EP model. With a much smaller model size and lower inference time, EP achieves near SotA ID performance and exhibits significantly improved OoD robustness compared to benchmark models. This work represents an initial step toward a trajectory prediction model that is capable of generalizing to different datasets and is robust under changes in sensor setup, scenario selection strategy or post-processing of training data. More sophisticated models based on this concept will be developed to improve the performance for both In- and Out-of-Distribution evaluations. Our work analyzes two important factors affecting model robustness: augmentation strategy and data representation. Other factors, such as model size, are not explored and remain to be researched.
@@ A-C Experiment Settings
TABLE V: Settings for EP Variants

### v2

@@ I Introduction
For practical application, the performance of trajectory prediction should be independent of these shared biases in a single dataset and measured by Out-of-Distribution (OoD) testing, e.g., across different datasets. As an example, Figure 1 clearly visualizes the distribution shift caused by different map generation processes between A2 and WO. However, the efforts for cross-dataset evaluation are hampered by the in-homogeneous data format and task details among datasets and competitions. Notable distinctions are the different lengths of scenario, observation history, and prediction horizon in each dataset.
Figure 1: Kernel density plot of the maximum absolute curvature and length for $5000$ random lane segments in A2 and WO. Contours indicate the $20$-th, $40$-th, $60$-th, and $80$-th percentiles, respectively.
@@
Figure 2: Our proposed model architecture. Inputs: Agent histories and road geometry are both represented via polynomials. Outputs: The tracked initial states and predicted states are fused into one polynomial trajectory prediction, ensuring continuity of past observation and future prediction.
@@ V-D1 Augmentation Strategy
Though heterogeneous augmentation does marginally reduce FMAE’s prediction error for both ID and OoD testing, the relative and absolute increase in prediction error between ID and OoD samples is practically the same as in the non-augmented case, e.g., $+0.279$m $(61.5\%)$ vs. $+0.295$m $(69.7\%)$ in $\Delta minADE_{6}$.
@@
On the other hand, the robustness of QCNet significantly benefits from the homogeneous augmentation with reduced $\Delta minFDE_{1}$ from $+1.802$m $(71.5\%)$ to $+1.067$m $(46.4\%)$. Compared to EP-noAug, EP-F and EP-Q replicate the behavior observed from FMAE and QCNet, e.g., with only a marginal impact from heterogeneous augmentation and a significant impact from homogeneous augmentation.
@@ V-D2 Data Representation
Our EP variants exhibit improved robustness compared to benchmarks with sequence-based data independent of data augmentation strategy. For instance, EP-F outperforms FMAE by demonstrating lower prediction errors and a lower increase in prediction errors in OoD testing. Similarly, compared to QCNet, EP-Q also demonstrates improved robustness, even achieving a reduction in prediction error of $-0.022$m $(-1.9\%)$ in $\Delta minADE_{1}$.
@@ V-D3 ID vs OoD results
We also examine the impact of model size on OoD results and scale the baselines to a comparable size to EP. Details and results are reported in Appendix B. Compared to the scaled baselines, our model still demonstrates improved robustness and even outperforms the scaled QCNet with lower prediction errors in $minADE_{6}$ and $minFDE_{6}$ in OoD testing.
@@ VI Conclusion
Recent SotA trajectory prediction models have thoroughly optimized their In-Distribution (ID) performance and present outstanding results in test sets evaluated in individual prediction competition. However, better ID performance does not automatically guarantee higher Out-of-Distribution (OoD) performance. In fact, performance differences can and do reverse in OoD settings. Our proposed OoD testing protocol enables a fresh perspective for model evaluation in trajectory prediction that we hope the community will adopt. With our protocol, we demonstrate the robustness improvement from homogenous augmentation and prove the benefits of polynomial representation as employed in our EP model. With a much smaller model size and lower inference time, EP achieves near SotA ID performance and exhibits significantly improved OoD robustness compared to benchmark models. This work represents an initial step toward a trajectory prediction model that is capable of generalizing to different datasets and is robust under changes in sensor setup, scenario selection strategy or post-processing of training data. More sophisticated models based on this concept will be developed to improve the performance for both In- and Out-of-Distribution evaluations.
@@ A-C Experiment Settings
Appendix B Scaling of Baselines
B-A Scaling Details of Baselines
We scale FMAE and QCNet to a comparable size to EP, with 313k and 413k model parameters, respectively. Scaling details are reported in Tables VI and VII. Other hyper-parameters remain the same for training the scaled baselines.
TABLE VI: Scaling Details of FMAE
  parameter modification
  embed dim $128\rightarrow 48$
num heads $8\rightarrow 4$
embed dim $512\rightarrow 180$
in lane_embedding layers $256\rightarrow 90$
TABLE VII: Scaling Details of QCNet
  parameter modification
  hidden dim $128\rightarrow 32$
head dim $16\rightarrow 8$
num heads $4\rightarrow 3$
num agent layers $2\rightarrow 1$
num dec layers $2\rightarrow 1$
B-B OoD Testing Results with Scaled Baselines
Figure 9: The OoD testing results of scaled FMAE (313k), scaled QCNet (413k), EP and their variants. We indicate the absolute and relative difference in displacement error between ID and OoD results. Solid: The ID results by training on the homogenized A2 training set and testing on the homogenized A2 validation set. Transparent: The increased displacement error in OoD testing by training on the homogenized A2 training set and testing on the homogenized WO validation set.